# On the fractal dimension of non-Newtonian Hele-Shaw flow subject to Saffman-Taylor instability

### v1

@@ Conclusion
In this work, we discuss the fractal dimension, $D_{\rm f}$, of the patterns that form due to the Saffman-Taylor instability of Hele-Shaw flow. We find that the shear thinning property of the more viscous fluid leads to an overall decrease of $D_{\rm f}$, compared to its Newtonian counterpart. This result is obtained from large-scale simulations of our proposed model, which modifies the well-known approach based on the DLA algorithm for simulating unstable Hele-Shaw two-fluid flow. Such a result would be difficult to obtain either experimentally or via continuous PDE-based simulations, since extremely large patterns are needed to produce accurate and converged results.
@@
The manner in which $D_{\rm f}$ changes due to non-Newtonian behavior is nontrivial. First, for vanishing surface tension, non-Newtonian behavior is not relevant since our model reduces to classical DLA simulations. For very large surface tension, the difference in $D_{\rm f}$ between Newtonian and non-Newtonian simulations becomes small, which may be expected in hindsight since the local curvature effects are dominant. For intermediate surface tension values, however, we find consistently smaller values of $D_{\rm f}$ for non-Newtonian fluids, compared to Newtonian ones. We hope that our findings will encourage new research, both experimental and theoretical, that will make progress in identifying the general connection between fluid rheological properties and fractal dimension of the emerging patterns.
Supplementary Material:
On the fractal dimension of non-Newtonian Hele-Shaw flow subject to Saffman-Taylor instability
I Monte Carlo simulations: modified Vicsek algorithm
@@
The main text provides the basic description of the Monte Carlo-based model that we use for simulating aggregate growth. The algorithm we use can simulate both Newtonian and shear-thinning non-Newtonian Hele-Shaw flow. We will first discuss the Newtonian aspects at the core of this algorithm and delay a discussion of the non-Newtonian modifications until Sections II and III.
@@
The algorithm is based on an irreversible growth model known as diffusion limited aggregation (DLA) [22]; see also [4] for a concise review. Briefly, DLA is a random walk on a lattice that leads to the formation of an aggregate from a seed particle. To align this with the Hele-Shaw geometry, the seed particle is placed at the origin of a planar rectangular grid. Subsequent particles are initialized one at a time at infinity (from a relatively large distance in the practical implementation) and randomly walk until they come in contact with the seeded aggregate, at which point they stick irreversibly. This process is then repeated until an aggregate of desired size is formed.
@@
$p_{\rm Vicsek}(n_{\cap})=A\left(\frac{n_{\cap}}{n_{\rm{total}}}-n_{0}\right)+B,$ (S1)
@@
The second modification we make is to use a different metric to define the local neighborhood of the walker. Instead of the rectangular metric used by Vicsek, we use a circular one. In other words, the walker’s neighborhood is now a circle instead of a square. This further requires us to reinterpret the parameters $n_{0}$ and $n_{\rm total}$ defining Eq. (S1). In the original formulation, Vicsek motivates the choice of these parameters by the fact that $n_{0}=(l-1)/(2l)$ corresponds to the number of lattice points contained by a rectangle of size $l\times(l-1)/2$ divided by the total number of lattice points $n_{\rm total}=l^{2}$. In other words, $n_{0}$ is the number of lattice points that would be occupied if the walker were in contact with a flat interface. As an example, if $l=5$, then the neighborhood is a $5\times 5$ square with $n_{\rm total}=25$ and $n_{0}=4/10=10/25.$
@@
In our modified approach, instead of a rectangle, we use a circle of diameter $l$ centered at the walker position and once again count the number of occupied sites within this circle. Since this circle is defined on top of the rectangular geometry specified by the grid, finding a general formula that expresses $n_{0}$ in terms of $l$ is difficult. For example, for $l=5$, $n_{\rm total}=13$ and $n_{0}=4/13$. In the case of $l=17,$ a value we use in all simulations throughout our work, we find by direct counting that $n_{0}=89/{n_{\rm total}}$, with $n_{\rm total}=193$ here and in Eq. (1) in the main text. Figure S1 illustrates the typical effects of using the modified versus the original algorithm. We find that the modified approaches encourage azimuthal symmetry of the emerging pattern for larger values of the surface tension parameter $A$.
@@
Figure S1: A comparison of modifying the neighborhood used in Newtonian simulations. All four simulations have $B=0.4$ and $l=17.$ The top two panels have $A=4$ while the bottom two panels have $A=16.$ The left panels use a square neighborhood, while the right use a circular one. All plots are on $1600\times 1600$ grids.
II Discussion of simulation parameters
Equations (2), (3) and the accompanying description in the main text explain how we further modify Vicsek’s algorithm to account for non-Newtonian effects due to a shear-thinning displaced fluid. We choose the parameters $C$ and $s$ in Eq. (2) (main text) in a manner that leads to an appreciable change in the sticking probability. This is done to mimic typical shear-thinning rheology: larger velocity magnitude of the interface leads to larger sticking probability, modeling smaller viscosity. Different non-Newtonian fluids have different responses to shear, thus, we can think of different values of these parameters as modeling different fluids.
To aid in deciding appropriate parameter values to use in our implementation, we plot the maximum value of the non-Newtonian contribution to the sticking probability, $p_{\rm NN}$, recorded over successively larger periods (labeled by integer $k$) of newly attached walkers (the length of each period is the largest integer of $1.3^{k+1}-1.3^{k}$, with the data points in Fig. S2 corresponding to $k\in[16,45]$). The choice of $1.3$ and the range of $k$ is arbitrary and was chosen simply for illustrative purposes.
We see that, for the chosen values of parameters $C$ and $s$, the desired qualitative behavior is observed, namely, for small numbers of walkers while the aggregate is increasing rapidly in size, the non-Newtonian correction is large, providing a sizeable non-Newtonian effect. This mimics the large shear-thinning effect anticipated at early times in an experiment, where velocity and shear-rate are large. At later times however, the (now large) aggregate is growing more slowly; one would anticipate a low shear-rate in the analogous physical fluid experiment, and the non-Newtonian correction to the sticking probability is accordingly much smaller.
Table 1 summarizes the parameter values used in our simulations.
Figure S2: The maximum non-Newtonian probability probability $p_{\rm NN}$ of Eq. (2) in the main text, running over each period of increasing length shown by the data points. The Newtonian parameters used are $A=4$ and $B=0.4$. Panel (a) uses $C=0.3,$ $s=0.2$ while panel (b) uses $C=0.5,$ $s=0.15$.
@@ II Discussion of simulation parameters
$l$ Radius of walkers’s neighborhood (nbd) 17
@@
$\left(C,s\right)$ Parameters defining $p_{\rm NN}$ $(0,1)\times(0,1)$
@@
III Further discussion of the attachment rules modeling shear thinning behavior
@@
In the main text, we discuss the concept of ‘age’. Here, we discuss the necessary algorithm modification relevant when the walker is in contact with exactly three occupied lattice sites when it attaches. Depending on how the walker is attached, the average of either their horizontal or vertical neighbors’ ages are used (e.g., for such a horizontal attachment, the horizontal age of the newly-attached walker is inherited from the single neighbor in the aggregate in horizontal contact, while vertical age is obtained by averaging over the two vertical neighbors). Figures S3 and S4 illustrate the algorithm for a few typical cases.
@@
Figure S3: Illustration of attachment rules modeling shear thinning behavior. White squares represent occupied lattice sites, while black ones represent unoccupied sites. In all panels, age is written in the form $\left(T_{H},~{}T_{V}\right)$ discussed in the main text, and the $k=1$ seed particle located at the origin has an age of (1,1). Panels (a) and (b) show simple examples of the horizontal and vertical age evolution of the aggregate in one dimension. Panels (c) and (d) show more elaborate examples where both the horizontal and vertical ages of the aggregate must be accounted for.
@@
Figure S4: An example where a walker, $(k=6),$ wants to attach at a lattice site in contact with three occupied sites. The red arrow indicates that this walker intends to attach horizontally from the left. In this case, the vertical age $T_{V}$ is the average of the vertical ages resulting from the $k=3$ and $k=5$ occupied sites, found from $((6-1)+(6-3))/2=4$, while the horizontal age $T_{H}$, found from $6-1=5$, is updated solely from the $k=1$ site.
IV Calculations of fractal dimension
@@
We characterize the morphologies of the resulting aggregates by computing their fractal dimensions. Since the fractal dimension is the crucial measure that we use for the quantification of the emerging patterns, we describe the algorithms implemented in detail here: box-counting (Sec. IV.1) and correlation length (Sec. IV.2) based.
IV.1 Box-counting Dimension
@@
$D_{{\rm f,box}}:=\lim_{\epsilon\rightarrow 0}\frac{\log{N\left(\epsilon\right)}}{\log{1/\epsilon}},$ (S2)
@@
In practice, the limit specified by Eq. (S2) is difficult to compute. To approximate it, we consider a nonzero limiting value of $\epsilon$ in Eq. (S2), and multiply both sides by $\log\left(1/\epsilon\right)$ to obtain a rough estimate
$N\left(\epsilon\right)\approx c\epsilon^{-D_{{\rm f,box}}},$ (S3)
@@
$B_{\epsilon}\left(\vec{x}\right)=\left\{\vec{y}:\left\|\vec{x}-\vec{y}\right\|_{\infty}\leq\epsilon\right\}.$ (S4)
@@
    Perform a linear fit of the logarithm of the number of covers versus logarithm of the size of the corresponding cover. The magnitude of the best-fit line’s slope is the box-counting dimension.
@@
IV.2 Correlation Length-based Dimension
The second approach we use to calculate the fractal dimension is through the correlation dimension [30]. First, we form the correlation integral $C(\epsilon)$ (adopting the nomenclature used by [31]), understood as the average probability of finding two particles within a ball of radius $\epsilon$, defined via a limit,
$C\left(\epsilon\right)=\lim_{N\rightarrow\infty}\frac{1}{N\left(N-1\right)}\sum_{\scriptstyle i,j=1\atop\scriptstyle i\neq j}^{N}\Theta\left(\epsilon-\left\|\vec{x}\left(i\right)-\vec{x}\left(j\right)\right\|_{l^{2}}\right),\quad\vec{x}\left(i\right)\in\mathbb{R}^{2},$ (S5)
@@ IV.2 Correlation Length-based Dimension
$D_{{\rm f,corr}}=\lim_{N\rightarrow\infty}\lim_{\epsilon\rightarrow 0^{+}}\frac{\log C\left(\epsilon\right)}{\log\epsilon}.$ (S6)
In practice, we estimate $C\left(\epsilon\right)$ by the finite correlation sum
$C_{N}\left(\epsilon\right)=\frac{1}{N\left(N-1\right)}\sum_{\scriptstyle i,j=1\atop\scriptstyle i\neq j}^{N}\Theta\left(\epsilon-\left\|\vec{x}\left(i\right)-\vec{x}\left(j\right)\right\|_{2}\right),\quad\vec{x}\left(i\right)\in\mathbb{R}^{2},$ (S7)
which is a viable approach due to the large number of particles we simulate, thus producing a very large $N$ and a reasonable approximation of Eq. (S5). Therefore, to estimate the correlation dimension using $C_{N}\left(\epsilon\right)$, we use
$D_{{\rm f,corr}}\approx\lim_{\epsilon\rightarrow 0^{+}}\frac{\log C_{N}\left(\epsilon\right)}{\log\epsilon}.$ (S8)
@@
A subtlety in the implementation of the correlation integral calculation is to avoid double counting in the double-indexed sum in Eq. (S7); double-counting arises from the fact that the intersection of $\epsilon$-neighborhoods of $\vec{x}_{i}$ and $\vec{x}_{j}$, $B_{\epsilon}\left(\vec{x}_{i}\right)\cap B_{\epsilon}\left(\vec{x}_{j}\right)$ is in general non-empty. To avoid double-counting, after we finish counting for $\vec{x}_{i}$, we remove this particle from the aggregate since we have counted all its unique pairs within its $\epsilon$-neighborhood (step 4 of Algorithm 2). A technicality of this implementation is that, since this removal procedure alters the aggregate, we must return to the original aggregate every time we change the search radius $\epsilon$.
V Additional Results
V.1 Reproducibility
In this section, we discuss the reproducibility of the fractal dimension results. To show that the results are realization independent, we set the parameters $A=1$ and $B=0.4$, and carry out two independent realizations of Newtonian and non-Newtonian simulations each (non-Newtonian parameters $C=0.9$ and $s=0.07$). Figure S5 shows the results of this numerical experiment for both fractal dimension calculation methods. The results, at sufficiently large aggregate sizes, are found to be consistent.
Figure S5: A consistency test for 2 sample simulations with Box Counting (circle $\circ$”) and Correlation dimension (“$\times$”). Panels (a) and (b) are Newtonian simulations (red). Panels (c) and (d) are non-Newtonian simulations (blue) with $C=0.9$ and $s=0.07$. In all cases, we use $A=1$ and $B=0.4$.
V.2 Influence of non-Newtonian parameters $(C,s)$
The parameters $(C,s)$ specify shear-thinning response, therefore modeling the rheology of the displaced fluid. Since we find that $D_{\rm f}$ differs between Newtonian and non-Newtonian simulations, it is not surprising that for different values of $(C,s)$ one could expect to find different values of $D_{\rm f}$. Table 3 shows that this is indeed the case; in addition, this table shows that that the change of $D_{\rm f}$, $\Delta D_{\rm f,\dagger}=D_{{\rm f,\dagger}}^{{\rm Newt}}-D_{{\rm f,\dagger}}^{{\rm nNewt}}$ (where $\dagger$ is ${\rm box}$ or ${\rm corr}$) is always non-negative, as pointed out also in the main text.

### v2

@@ V Conclusions
In this work, we discuss the fractal dimension, $D_{\rm f}$, of the patterns that form due to the Saffman-Taylor instability of Hele-Shaw flow. We find that the shear thinning property of the more viscous fluid leads to an overall decrease of $D_{\rm f}$, compared to its Newtonian counterpart. This result is obtained from large-scale simulations of our proposed model, which modifies the well-known approach based on the DLA algorithm for simulating unstable Hele-Shaw two-fluid flow. Such a result would be difficult to obtain either experimentally or via continuous PDE-based simulations since extremely large patterns are needed to produce accurate and converged results.
@@
The manner in which $D_{\rm f}$ changes due to non-Newtonian behavior is nontrivial. First, for vanishing surface tension, non-Newtonian behavior is not relevant since our model reduces to classical DLA simulations. For very large surface tension, the difference in $D_{\rm f}$ between Newtonian and non-Newtonian simulations becomes small, which may be expected in hindsight since the local curvature effects are dominant. For intermediate surface tension values, however, we find consistently smaller values of $D_{\rm f}$ for non-Newtonian fluids, compared to Newtonian ones. We hope that our findings will encourage new research, both experimental and theoretical, that will make progress in identifying the general connection between fluid rheological properties and the fractal dimension of the emerging patterns.
Acknowledgements.
We hereby acknowledge three cohorts of NJIT undergraduate students (too many to list) who worked on experiments and preliminary versions of the model presented here, with the last iteration leading to the results given in [8].
Appendix A Monte Carlo simulations: modified Vicsek algorithm
@@
Section III provides the basic description of the Monte Carlo-based model that we use for simulating aggregate growth. The algorithm we use can simulate both Newtonian and shear-thinning non-Newtonian Hele-Shaw flow. We will first discuss the Newtonian aspects at the core of this algorithm and delay a discussion of the non-Newtonian modifications until Appendices B and C.
@@
The algorithm is based on an irreversible growth model known as diffusion-limited aggregation (DLA) [22]; see also [4] for a concise review. Briefly, DLA is a random walk on a lattice that leads to the formation of an aggregate from a seed particle. To align this with the Hele-Shaw geometry, the seed particle is placed at the origin of a planar rectangular grid. Subsequent particles are initialized one at a time at infinity (from a relatively large distance in the practical implementation) and randomly walk until they come in contact with the seeded aggregate, at which point they stick irreversibly. This process is then repeated until an aggregate of the desired size is formed.
@@
$p_{\rm Vicsek}(n_{\cap})=A\left(\frac{n_{\cap}}{n_{\rm{total}}}-n_{0}\right)+B,$ (4)
@@
The second modification we make is to use a different metric to define the local neighborhood of the walker. Instead of the rectangular metric used by Vicsek, we use a circular one. In other words, the walker’s neighborhood is now a circle instead of a square. This further requires us to reinterpret the parameters $n_{0}$ and $n_{\rm total}$ defining Eq. (4). In the original formulation, Vicsek motivates the choice of these parameters by the fact that $n_{0}=(l-1)/(2l)$ corresponds to the number of lattice points contained by a rectangle of size $l\times(l-1)/2$ divided by the total number of lattice points $n_{\rm total}=l^{2}$. In other words, $n_{0}$ is the number of lattice points that would be occupied if the walker were in contact with a flat interface. As an example, if $l=5$, then the neighborhood is a $5\times 5$ square with $n_{\rm total}=25$ and $n_{0}=4/10=10/25.$
@@
In our modified approach, instead of a rectangle, we use a circle of diameter $l$ centered at the walker position and once again count the number of occupied sites within this circle. Since this circle is defined on top of the rectangular geometry specified by the grid, finding a general formula that expresses $n_{0}$ in terms of $l$ is difficult. For example, for $l=5$, $n_{\rm total}=13$ and $n_{0}=4/13$. In the case of $l=17,$ a value we use in all simulations throughout our work, we find by direct counting that $n_{0}=89/{n_{\rm total}}$, with $n_{\rm total}=193$ here and in Eq. (1). Figure 4 illustrates the typical effects of using the modified versus the original algorithm. We find that the modified approaches encourage azimuthal symmetry of the emerging pattern for larger values of the surface tension parameter $A$.
@@
Figure 4: An influence of modifying the neighborhood used in Newtonian simulations. All four simulations have $B=0.4$ and $l=17.$ The top two panels have $A=4$ while the bottom two panels have $A=16.$ The left panels use a square neighborhood, while the right uses a circular one. All plots are on $1600\times 1600$ grids.
Appendix B Discussion of simulation parameters
@@
$l$ Radius of walkers’ neighborhood (nbd) 17
@@
$\left(C,s\right)$ Parameters defining $p_{\rm nNewt}$ $(0,1)\times(0,1)$
Equations (2), (3) and the accompanying description explain how we further modify Vicsek’s algorithm to account for non-Newtonian effects due to a shear-thinning displaced fluid. We choose the parameters $C$ and $s$ in Eq. (2) in a manner that leads to an appreciable change in the sticking probability. This is done to mimic typical shear-thinning rheology: larger velocity magnitude of the interface leads to a larger sticking probability, modeling smaller viscosity. Different non-Newtonian fluids have different responses to shear, thus, we can think of different values of these parameters as modeling different fluids.
To aid in deciding appropriate parameter values to use in our implementation, we plot the maximum value of the non-Newtonian contribution to the sticking probability, $p_{\rm nNewt}$, recorded over successively larger periods (labeled by integer $k$) of newly attached walkers (the length of each period is the largest integer of $1.3^{k+1}-1.3^{k}$, with the data points in Fig. 5 corresponding to $k\in[16,45]$). The choice of $1.3$ and the range of $k$ is arbitrary and was chosen simply for illustrative purposes.
We see that, for the chosen values of parameters $C$ and $s$, the desired qualitative behavior is observed, namely, for small numbers of walkers while the aggregate is increasing rapidly in size, the non-Newtonian correction is large, providing a sizeable non-Newtonian effect. This mimics the large shear-thinning effect anticipated at early times in an experiment, where velocity and shear rate are large. At later times, however, the (now large) aggregate is growing more slowly; one would anticipate a low shear rate in the analogous physical fluid experiment, and the non-Newtonian correction to the sticking probability is accordingly much smaller.
Table 1 summarizes the parameter values used in our simulations.
Figure 5: The maximum non-Newtonian probability probability $p_{\rm nNewt}$ of Eq. (2), running over each period of increasing length shown by the data points. The Newtonian parameters used are $A=4$ and $B=0.4$. Panel (a) uses $C=0.3,$ $s=0.2$ while panel (b) uses $C=0.5,$ $s=0.15$.
@@
Appendix C Further discussion of the attachment rules modeling shear thinning behavior
@@
In Sec. III, we discuss the concept of ‘age’. Here, we discuss the necessary algorithm modification relevant when the walker is in contact with exactly three occupied lattice sites when it attaches. Depending on how the walker is attached, the average of either their horizontal or vertical neighbors’ ages are used (e.g., for such a horizontal attachment, the horizontal age of the newly-attached walker is inherited from the single neighbor in the aggregate in horizontal contact, while vertical age is obtained by averaging over the two vertical neighbors). Figures 6 and 7 illustrate the algorithm for a few typical cases.
@@
Figure 6: Illustration of attachment rules modeling shear thinning behavior. White squares represent occupied lattice sites, while black ones represent unoccupied sites. In all panels, age is written in the form $\left(T_{H},~{}T_{V}\right)$, and the $k=1$ seed particle located at the origin has an age of (1,1). Panels (a) and (b) show simple examples of the horizontal and vertical age evolution of the aggregate in one dimension. Panels (c) and (d) show more elaborate examples where both the horizontal and vertical ages of the aggregate must be accounted for.
@@
Figure 7: An example where a walker, $(k=6),$ intends to attach at a lattice site in contact with three occupied sites. The red arrow indicates that this walker intends to attach horizontally from the left. In this case, the vertical age $T_{V}$ is the average of the vertical ages resulting from the $k=3$ and $k=5$ occupied sites, found from $((6-1)+(6-3))/2=4$, while the horizontal age $T_{H}$, found from $6-1=5$, is updated solely from the $k=1$ site.
Appendix D Calculations of fractal dimension
@@
We characterize the morphologies of the resulting aggregates by computing their fractal dimensions. Since the fractal dimension is the crucial measure that we use for the quantification of the emerging patterns, we describe the algorithms implemented in detail here, based on box-counting (Appendix D.1) and correlation length (Appendix D.2).
D.1 Box-counting Dimension
@@
$D_{{\rm f,box}}:=\lim_{\epsilon\rightarrow 0}\frac{\log{N\left(\epsilon\right)}}{\log{1/\epsilon}},$ (5)
@@
In practice, the limit specified by Eq. (5) is difficult to compute. To approximate it, we consider a nonzero limiting value of $\epsilon$ in Eq. (5), and multiply both sides by $\log\left(1/\epsilon\right)$ to obtain a rough estimate
$N\left(\epsilon\right)\approx c\epsilon^{-D_{{\rm f,box}}},$ (6)
@@
$B_{\epsilon}\left(\vec{x}\right)=\left\{\vec{y}:\left\|\vec{x}-\vec{y}\right\|_{\infty}\leq\epsilon\right\}.$ (7)
@@
    Perform a linear fit of the logarithm of the number of covers versus the logarithm of the size of the corresponding cover. The magnitude of the best-fit line’s slope is the box-counting dimension.
@@
D.2 Correlation Length-based Dimension
The second approach we use to calculate the fractal dimension is through the correlation dimension [29]. First, we form the correlation integral $C(\epsilon)$ (adopting the nomenclature used by [30]), understood as the average probability of finding two particles within a ball of radius $\epsilon$, defined via a limit, for $\vec{x}\left(i\right)\in\mathbb{R}^{2}$,
$C\left(\epsilon\right)=\lim_{N\rightarrow\infty}\frac{1}{N\left(N-1\right)}\sum_{\scriptstyle i,j=1\atop\scriptstyle i\neq j}^{N}\Theta\left(\epsilon-\left\|\vec{x}\left(i\right)-\vec{x}\left(j\right)\right\|_{l^{2}}\right),$ (8)
@@ D.2 Correlation Length-based Dimension
$D_{{\rm f,corr}}=\lim_{N\rightarrow\infty}\lim_{\epsilon\rightarrow 0^{+}}\frac{\log C\left(\epsilon\right)}{\log\epsilon}.$ (9)
In practice, we estimate $C\left(\epsilon\right)$ by the finite correlation sum, for $\vec{x}\left(i\right)\in\mathbb{R}^{2}$,
$C_{N}\left(\epsilon\right)=\frac{1}{N\left(N-1\right)}\sum_{\scriptstyle i,j=1\atop\scriptstyle i\neq j}^{N}\Theta\left(\epsilon-\left\|\vec{x}\left(i\right)-\vec{x}\left(j\right)\right\|_{2}\right),$ (10)
which is a viable approach due to the large number of particles we simulate, thus producing a very large $N$ and a reasonable approximation of Eq. (8). Therefore, to estimate the correlation dimension using $C_{N}\left(\epsilon\right)$, we use
$D_{{\rm f,corr}}\approx\lim_{\epsilon\rightarrow 0^{+}}\frac{\log C_{N}\left(\epsilon\right)}{\log\epsilon}.$ (11)
@@
A subtlety in the implementation of the correlation integral calculation is to avoid double counting in the double-indexed sum in Eq. (10); double-counting arises from the fact that the intersection of $\epsilon$-neighborhoods of $\vec{x}_{i}$ and $\vec{x}_{j}$, $B_{\epsilon}\left(\vec{x}_{i}\right)\cap B_{\epsilon}\left(\vec{x}_{j}\right)$ is in general non-empty. To avoid double-counting, after we finish counting for $\vec{x}_{i}$, we remove this particle from the aggregate since we have counted all its unique pairs within its $\epsilon$-neighborhood (step 4 of Algorithm 2). A technicality of this implementation is that, since this removal procedure alters the aggregate, we must return to the original aggregate every time we change the search radius $\epsilon$.
Appendix E Additional Results
E.1 Reproducibility
Figure 8: A consistency test for 2 sample simulations with Box Counting (circle $\circ$”) and Correlation dimension (“$\times$”). Panels (a) and (b) are Newtonian simulations (red). Panels (c) and (d) are non-Newtonian simulations (blue) with $C=0.9$ and $s=0.07$. In all cases, we use $A=1$ and $B=0.4$.
In this Appendix, we discuss the reproducibility of the fractal dimension results. To show that the results are realization independent, we set the parameters $A=1$ and $B=0.4$, and carry out two independent realizations of Newtonian and non-Newtonian simulations each (non-Newtonian parameters $C=0.9$ and $s=0.07$). Figure 8 shows the results of this numerical experiment for both fractal dimension calculation methods. The results, at sufficiently large aggregate sizes, are found to be consistent.
E.2 Influence of non-Newtonian parameters $(C,s)$
The parameters $(C,s)$ specify shear-thinning response, therefore modeling the rheology of the displaced fluid. Since we find that $D_{\rm f}$ differs between Newtonian and non-Newtonian simulations, it is not surprising that for different values of $(C,s)$ one could expect to find different values of $D_{\rm f}$. Table 3 shows that this is indeed the case; in addition, this table shows that that the change of $D_{\rm f}$, $\Delta D_{\rm f,\dagger}=D_{{\rm f,\dagger}}^{{\rm Newt}}-D_{{\rm f,\dagger}}^{{\rm nNewt}}$ (where $\dagger$ is ${\rm box}$ or ${\rm corr}$) is always non-negative.